OpenStudy (chupacabraj):

HELP An electron is launched horizontally to the right with an initial speed of 100 m/s in a uniform horizontal electric field of 10,000 v/m where field lines point from left to right. How long after launch until the electron's speed is zero?

OpenStudy (chupacabraj):

f=ma, so use F=Eq

OpenStudy (chupacabraj):

for f get 1.6 x 10^-15 N

OpenStudy (chupacabraj):

Using kinematics to solve for t I get a negative, and the absolute value of that isn't even the solution

OpenStudy (irishboy123):

\(a = \dfrac{10 \cdot 10^3 \times -1.6 \cdot 10^{-19}}{9.1 \cdot 10^{-31}} = -1.75 \cdot 10^{15}\) for the electron, \(KE_o = 0.5 \cdot 10^{-26}\) and the energy radiated away during the deceleration is \(\approx 10^{-46} \)J, so that's insignificant. IOW (non-relativistic) kinematics should give the answer

OpenStudy (chupacabraj):

Ohh I forgot about the negative sign since it's an electron!

OpenStudy (chupacabraj):

Vf = Vi + at?

OpenStudy (chupacabraj):

Is that the correct kinematic equation to use in order to get the time at which v= 0?

OpenStudy (chupacabraj):

This is the answer

OpenStudy (irishboy123):

yes \(v_f = u_i + at\) \(t = \dfrac{0-100}{−1.75⋅10^{15}} \approx 5.7 \cdot 10^{-14}\) check the physical situation first to avoid nasty surprises. if the E field is left to right, a proton (+ve charge) will move left to right in that field. that's how the E field vector is defined. and so an electron will want to move right to left getting the signs in the formula right is then easier the answer book looks wrong :(

OpenStudy (chupacabraj):

It looks similiar, perhaps they decided to keep it in fraction form? I just wonder how they'd do it though. They won't let me use a calculator. :S